OpenStudy (anonymous):

prove: csc x+cot x=sinx/1-cosx

OpenStudy (danjs):

\[\frac{ 1 }{ \sin(x) } + \frac{ \cos(x) }{ \sin(x) } = \frac{ \sin(x) }{ 1-\cos(x) }\]

OpenStudy (danjs):

is that the correct right side? you dont have a parenthesis around the (1-cos(x))

OpenStudy (anonymous):

theres no parenthesis at all

OpenStudy (danjs):

then the right side is just \[\frac{ \sin(x) }{ 1 } - \cos(x)\] ?

OpenStudy (anonymous):

you we right at first but theres no parenthesis at all including both sides

Nnesha (nnesha):

do you mean this  like this ??? :D